# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ Appendix E Consistency under Diverging Candidate Cardinality
$\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)\lesssim\exp(-K^{\prime}(\log^{-1}n\cdot n^{\mathcal{C}_{n}/2})^{\theta^{\prime}}).$ (73)
@@
pr( correct selection )$\operatorname{pr}(\text{ correct selection })$ (74)
@@
$\quad-\operatorname{pr}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$
@@
Establishing the last equality above is identical as in the proof of Theorem 4.1, leveraging assumption (U2) that all the alternative models are worse than $\{g_{i}\}$. So long as
@@ E.1 Technical Results for Exponential Concentration
    martingale property: $E\left(D_{i}\mid D_{j}:j\in[i-1]\right)=0$ for all $2\leq i\leq n$, and $ED_{1}=0$.
@@
$\|D\|_{q}\leq c^{\prime}\left(\sum_{i=1}^{n}K_{i}^{2}\right)^{1/2}q^{1/\alpha^{\prime}},\quad\forall q\geq 1.$ (77)
@@
$\|D\|_{q}\leq c^{\prime}\sqrt{n}Kq^{1/\alpha^{\prime}},\quad\forall q\geq 1.$ (78)
@@
$\|D\|_{q}\leq\left[(q-1)\sum_{i=1}^{n}\left\|D_{i}\right\|_{q}^{2}\right]^{1/2}\leq\left[C(q-1)q^{2/\alpha^{*}}\sum_{i=1}^{n}K_{i}^{2}\right]^{1/2}\leq C^{1/2}q^{\frac{2+\alpha^{*}}{2\alpha^{*}}}\left(\sum_{i=1}^{n}K_{i}^{2}\right)^{1/2}$ (79)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}^{(k)}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)\leq\exp\left(-K^{\prime}\left(c_{n}/\mathbb{c}_{n}^{(k)}\right)^{\theta^{\prime}}\right),$ (80)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}^{(k)}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)\leq\exp\left(-K^{\prime}n^{\theta^{\prime}(1-a_{k})/2}\right).$ (81)
@@
$E\left[D_{i}\mid D_{j}:j\in[i-1]\right]=0$
for all $2\leq i\leq n$, and $ED_{1}=0$ ($\epsilon_{i}$ is centered conditioning on $X_{i}$).
@@
$E[(D_{i})^{q}]$ (82)
$=i^{\xi q}E\left[\left(\epsilon_{i}\right)^{q}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{q}\right]$
$\leq i^{\xi q}i^{-aq/2}\sqrt{E\left[\epsilon_{i}^{2q}\right]}\sqrt{E\left[i^{aq}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{2q}\right]}$
@@
$\Rightarrow(E[(D_{i})^{q}])^{1/q}$ $\leq i^{\xi-a/2}c^{2}K_{\epsilon}K_{\rm est}q^{\theta_{\epsilon}^{-1}+\theta_{\rm est}^{-1}}$
In step (I) we used the moment bounds for sub-Weilbull random variables. So $D_{i}$ is $(ci^{\xi-a/2}K_{\epsilon}K_{\rm est},(\theta_{\epsilon}^{-1}+\theta_{\rm est}^{-1})^{-1})$-sub-Weibull. Given the martingale structure and sub-Weibull conditions for each $D_{i}$, we can employ Lemma E.1 to derive sub-Weibull properties for $\sum_{i=1}^{n}D_{i}$:
$\left\|\sum_{i=1}^{n}D_{i}\right\|_{q}$ $\leq c\left(\sum_{i=1}^{n}\left(K_{\epsilon}K_{\rm est}\right)^{2}i^{2\xi-a}\right)^{1/2}q^{1/\theta^{\prime}}\text{ for all }q\geq 1$ (83)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)$ (84)
$=\operatorname{pr}\left(\frac{\left|\sum_{i=1}^{n}D_{i}\right|}{K_{\epsilon}K_{\rm est}\mathbb{c}_{n}}\geq\frac{c_{n}}{K_{\epsilon}K_{\rm est}\mathbb{c}_{n}}\right)$
@@
We can similarly derive a bound for $\sum_{i=1}^{n}u_{i}^{(k)}-E\left[u_{i}^{(k)}\mid F^{i-1}\right]$.
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}u_{i}^{(k)}-E\left[u_{i}^{(k)}\mid F^{i-1}\right]\right|\geq c_{n}\right)\leq\exp\left(-K^{\prime}\left(c_{n}n^{-\xi+a_{k}-1/2}\right)^{\theta^{\prime}}\right)$ (85)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}u_{i}^{(k)}-E\left[u_{i}^{(k)}\mid F^{i-1}\right]\right|\geq c_{n}\right)\leq\exp(-K^{\prime}n^{\theta^{\prime}/2}).$ (86)
@@
We drop the $k$ in the superscript for notation simplicity. For each $k$, denote $D_{i}=u_{i}-E\left[u_{i}\mid F^{i-1}\right]$, we need to derive some moments bounds on $D_{i}$.
$\|D_{i}\|_{q}$ $\leq\|u_{i}\|_{q}+\|E[u_{i}\mid F^{i-1}]\|_{q}$ (87)
@@
$=2\left(E\left[\left(i^{\xi}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{2}\right)^{q}\right]\right)^{1/q}$
@@
$\left\|\sum_{i=1}^{n}D_{i}\right\|_{q}$ $\leq c\left(\sum_{i=1}^{n}C\left(K_{\text{est }},\theta_{\mathrm{est}}\right)i^{2\xi-2a}\right)^{1/2}q^{2/\theta_{\text{est }}}$ (88)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}u_{i}-E\left[u_{i}\mid F^{i-1}\right]\right|\geq c_{n}\right)$ (89)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}E\left[u_{i}^{(k)}\mid F^{i-1}\right]-E\left[u_{i}^{(k)}\right]\right|\geq c_{n}\right)\leq\exp\left[-K^{\prime}\left\{c_{n}\left(n^{-\mathbb{c}-3/2}\log^{-1}n\wedge\log^{-1}n\right)\right\}^{\theta^{\prime}}\right],$ (90)
@@
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}E\left[u_{i}^{(k)}\mid F^{i-1}\right]-E\left[u_{i}^{(k)}\right]\right|\geq c_{n}\right)\leq\exp\left[-K^{\prime}\left\{\log^{-1}n\left(n^{b_{k}-a_{k}/2-1/2}\wedge n^{1-a_{k}+\xi}\right)\right\}^{\theta^{\prime}}\right].$ (91)
@@
$\sum_{i=1}^{n}E\left[u_{i}\mid F^{i-1}\right]-\sum_{i=1}^{n}E\left[u_{i}\right]$ (92)
$=$ $\sum_{i=1}^{n}\left(E\left[u_{i}\mid F^{i-1}\right]-E\left[u_{i}\mid F^{i-2}\right]+E\left[u_{i}\mid F^{i-2}\right]-E\left[u_{i}\mid F^{i-3}\right]+\cdots+E\left[u_{i}\mid F^{1}\right]-E\left[u_{i}\right]\right)$
$=$ $\sum_{i=1}^{n}\sum_{j=1}^{i-1}E\left[u_{i}\mid F^{j}\right]-E\left[u_{i}\mid F^{j-1}\right]$
$=$ $\sum_{j=1}^{n-1}\sum_{i=j+1}^{n}E\left[u_{i}\mid F^{j}\right]-E\left[u_{i}\mid F^{j-1}\right]=:\sum_{j=1}^{n-1}D_{j},$
where $D_{j}:=\sum_{i=j+1}^{n}E\left[u_{i}\mid F^{j}\right]-E\left[u_{i}\mid F^{j-1}\right]$. Note that $D_{j}$ is a martingale difference sequence
$E[D_{j}\mid D_{k}:k\in[j-1]]=E[E\{D_{j}\mid D_{k}:k\in[j-1],F^{j-1}\}]=0.$
@@
$=$ $\left\|\sum_{i=j+1}^{n}E\left[u_{i}\mid F^{j}\right]-E\left[u_{i}\mid F^{j-1}\right]\right\|_{q}$
$\leq$ $\sum_{i=j+1}^{n}\left\|E\left[u_{i}\mid F^{j}\right]-E\left[u_{i}\mid F^{j-1}\right]\right\|_{q}$
$=$ $\sum_{i=j+1}^{n}\left\|E\left[\nabla_{j}u_{i}\mid F^{j}\right]\right\|_{q}$
@@
$\nabla_{j}u_{i}=i^{\xi}\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)+\hat{f}_{i-1}\left(X_{i};Z_{i-1}^{j}\right)-f_{0}\left(X_{i}\right)\right\}.$ (94)
@@
$\left\|\nabla_{j}u_{i}\right\|_{q}$ (95)
@@
$\left\|i^{\xi}\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)\right\}\right\|_{q}^{q}$ (96)
$\leq i^{\xi q}\sqrt{E\left[\left(\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right)^{2q}\right]}\sqrt{E\left[\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)\right\}^{2q}\right]}$
$=i^{(\xi-b-a/2)q}\sqrt{E\left[i^{2qb}\left(\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right)^{2q}\right]}\sqrt{E\left[i^{aq}\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)\right\}^{2q}\right]}$
@@
$\mathbb{c}_{n}=\sqrt{\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}i^{\mathbb{c}}\right)^{2}}\leq c(n^{3/2+\mathbb{c}}\log n\vee\log n).$ (97)
So overall we have
$\operatorname{pr}\left(\left|\sum_{i=1}^{n}E\left[u_{i}\mid F^{i-1}\right]-E\left[u_{i}\right]\right|\geq c_{n}\right)$ (98)
$=$ $\operatorname{pr}\left(\left|\sum_{i=1}^{n}E\left[u_{i}\mid F^{i-1}\right]-E\left[u_{i}\right]\right|/\mathbb{c}_{n}\geq c_{n}/\mathbb{c}_{n}\right)$

### v4

@@ Appendix E Consistency under Diverging Candidate Cardinality
$\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{h_{i-1}^{(k)}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\leq(1-\delta)H_{n}\right)\lesssim\exp(-K^{\prime}(\log^{-1}n\cdot n^{\mathcal{C}_{n}/2})^{\theta^{\prime}}).$
@@
( correct selection ) $\mathbb{P}(\text{ correct selection })$
@@
$\quad-\mathbb{P}\left(\sum_{i=1}^{n}i^{\xi}\left\{g_{i-1}\left(X_{i}\right)-Y_{i}\right\}^{2}-i^{\xi}\epsilon_{i}^{2}\geq(1-\delta)H_{n}\right)$
@@
Establishing the last equality above is identical to the proof of Theorem 4.1, leveraging assumption (U2) that all the alternative models are worse than $\{g_{i}\}$. So long as
@@ E.1 Technical Results for Exponential Concentration
    martingale property: $\mathbb{E}\left(D_{i}\mid D_{j}:j\in[i-1]\right)=0$ for all $2\leq i\leq n$, and $\mathbb{E}D_{1}=0$.
@@
$\|D\|_{q}\leq c^{\prime}\left(\sum_{i=1}^{n}K_{i}^{2}\right)^{1/2}q^{1/\alpha^{\prime}},\quad\forall q\geq 1.$
@@
$\|D\|_{q}\leq c^{\prime}\sqrt{n}Kq^{1/\alpha^{\prime}},\quad\forall q\geq 1.$
@@
$\|D\|_{q}\leq\left[(q-1)\sum_{i=1}^{n}\left\|D_{i}\right\|_{q}^{2}\right]^{1/2}\leq\left[C(q-1)q^{2/\alpha^{*}}\sum_{i=1}^{n}K_{i}^{2}\right]^{1/2}\leq C^{1/2}q^{\frac{2+\alpha^{*}}{2\alpha^{*}}}\left(\sum_{i=1}^{n}K_{i}^{2}\right)^{1/2}$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}^{(k)}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)\leq\exp\left(-K^{\prime}\left(c_{n}/\mathbb{c}_{n}^{(k)}\right)^{\theta^{\prime}}\right),$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}^{(k)}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)\leq\exp\left(-K^{\prime}n^{\theta^{\prime}(1-a_{k})/2}\right).$
@@
$\mathbb{E}\left[D_{i}\mid D_{j}:j\in[i-1]\right]=0$
for all $2\leq i\leq n$, and $\mathbb{E}D_{1}=0$ ($\epsilon_{i}$ is centered conditioning on $X_{i}$).
@@
$\mathbb{E}[(D_{i})^{q}]$
$=i^{\xi q}\mathbb{E}\left[\left(\epsilon_{i}\right)^{q}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{q}\right]$
$\leq i^{\xi q}i^{-aq/2}\sqrt{\mathbb{E}\left[\epsilon_{i}^{2q}\right]}\sqrt{\mathbb{E}\left[i^{aq}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{2q}\right]}$
@@
$\Rightarrow(\mathbb{E}[(D_{i})^{q}])^{1/q}$ $\leq i^{\xi-a/2}c^{2}K_{\epsilon}K_{\rm est}q^{\theta_{\epsilon}^{-1}+\theta_{\rm est}^{-1}}$
In step (I), we used the moment bounds for sub-Weibull random variables. So $D_{i}$ is $(ci^{\xi-a/2}K_{\epsilon}K_{\rm est},(\theta_{\epsilon}^{-1}+\theta_{\rm est}^{-1})^{-1})$-sub-Weibull. Given the martingale structure and sub-Weibull conditions for each $D_{i}$, we can employ Lemma E.1 to derive sub-Weibull properties for $\sum_{i=1}^{n}D_{i}$:
$\left\|\sum_{i=1}^{n}D_{i}\right\|_{q}$ $\leq c\left(\sum_{i=1}^{n}\left(K_{\epsilon}K_{\rm est}\right)^{2}i^{2\xi-a}\right)^{1/2}q^{1/\theta^{\prime}}\text{ for all }q\geq 1$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}i^{\xi}\epsilon_{i}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}\right|\geq c_{n}\right)$
$=\mathbb{P}\left(\frac{\left|\sum_{i=1}^{n}D_{i}\right|}{K_{\epsilon}K_{\rm est}\mathbb{c}_{n}}\geq\frac{c_{n}}{K_{\epsilon}K_{\rm est}\mathbb{c}_{n}}\right)$
@@
We can similarly derive a bound for $\sum_{i=1}^{n}u_{i}^{(k)}-\mathbb{E}\left[u_{i}^{(k)}\mid F^{i-1}\right]$.
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}u_{i}^{(k)}-\mathbb{E}\left[u_{i}^{(k)}\mid F^{i-1}\right]\right|\geq c_{n}\right)\leq\exp\left(-K^{\prime}\left(c_{n}n^{-\xi+a_{k}-1/2}\right)^{\theta^{\prime}}\right)$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}u_{i}^{(k)}-\mathbb{E}\left[u_{i}^{(k)}\mid F^{i-1}\right]\right|\geq c_{n}\right)\leq\exp(-K^{\prime}n^{\theta^{\prime}/2}).$
@@
We drop the $k$ in the superscript for notation simplicity. For each $k$, denote $D_{i}=u_{i}-\mathbb{E}\left[u_{i}\mid F^{i-1}\right]$, we need to derive some moments bounds on $D_{i}$.
$\|D_{i}\|_{q}$ $\leq\|u_{i}\|_{q}+\|\mathbb{E}[u_{i}\mid F^{i-1}]\|_{q}$
@@
$=2\left(\mathbb{E}\left[\left(i^{\xi}\left\{\hat{f}_{i-1}\left(X_{i}\right)-f_{0}\left(X_{i}\right)\right\}^{2}\right)^{q}\right]\right)^{1/q}$
@@
$\left\|\sum_{i=1}^{n}D_{i}\right\|_{q}$ $\leq c\left(\sum_{i=1}^{n}C\left(K_{\text{est }},\theta_{\mathrm{est}}\right)i^{2\xi-2a}\right)^{1/2}q^{2/\theta_{\text{est }}}$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}u_{i}-\mathbb{E}\left[u_{i}\mid F^{i-1}\right]\right|\geq c_{n}\right)$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}\mathbb{E}\left[u_{i}^{(k)}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}^{(k)}\right]\right|\geq c_{n}\right)\leq\exp\left[-K^{\prime}\left\{c_{n}\left(n^{-\mathbb{c}-3/2}\log^{-1}n\wedge\log^{-1}n\right)\right\}^{\theta^{\prime}}\right],$
@@
$\mathbb{P}\left(\left|\sum_{i=1}^{n}\mathbb{E}\left[u_{i}^{(k)}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}^{(k)}\right]\right|\geq c_{n}\right)\leq\exp\left[-K^{\prime}\left\{\log^{-1}n\left(n^{b_{k}-a_{k}/2-1/2}\wedge n^{1-a_{k}+\xi}\right)\right\}^{\theta^{\prime}}\right].$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[u_{i}\mid F^{i-1}\right]-\sum_{i=1}^{n}\mathbb{E}\left[u_{i}\right]$
$=$ $\sum_{i=1}^{n}\left(\mathbb{E}\left[u_{i}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}\mid F^{i-2}\right]+\mathbb{E}\left[u_{i}\mid F^{i-2}\right]-\mathbb{E}\left[u_{i}\mid F^{i-3}\right]+\cdots+\mathbb{E}\left[u_{i}\mid F^{1}\right]-\mathbb{E}\left[u_{i}\right]\right)$
$=$ $\sum_{i=1}^{n}\sum_{j=1}^{i-1}\mathbb{E}\left[u_{i}\mid F^{j}\right]-\mathbb{E}\left[u_{i}\mid F^{j-1}\right]$
$=$ $\sum_{j=1}^{n-1}\sum_{i=j+1}^{n}\mathbb{E}\left[u_{i}\mid F^{j}\right]-\mathbb{E}\left[u_{i}\mid F^{j-1}\right]=:\sum_{j=1}^{n-1}D_{j},$
where $D_{j}:=\sum_{i=j+1}^{n}\mathbb{E}\left[u_{i}\mid F^{j}\right]-\mathbb{E}\left[u_{i}\mid F^{j-1}\right]$. Note that $D_{j}$ is a martingale difference sequence
$\mathbb{E}[D_{j}\mid D_{k}:k\in[j-1]]=\mathbb{E}[E\{D_{j}\mid D_{k}:k\in[j-1],F^{j-1}\}]=0.$
@@
$=$ $\left\|\sum_{i=j+1}^{n}\mathbb{E}\left[u_{i}\mid F^{j}\right]-\mathbb{E}\left[u_{i}\mid F^{j-1}\right]\right\|_{q}$
$\leq$ $\sum_{i=j+1}^{n}\left\|\mathbb{E}\left[u_{i}\mid F^{j}\right]-\mathbb{E}\left[u_{i}\mid F^{j-1}\right]\right\|_{q}$
$=$ $\sum_{i=j+1}^{n}\left\|\mathbb{E}\left[\nabla_{j}u_{i}\mid F^{j}\right]\right\|_{q}$
@@
$\nabla_{j}u_{i}=i^{\xi}\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)+\hat{f}_{i-1}\left(X_{i};Z_{i-1}^{j}\right)-f_{0}\left(X_{i}\right)\right\}.$
@@
$\left\|\nabla_{j}u_{i}\right\|_{q}$
@@
$\left\|i^{\xi}\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)\right\}\right\|_{q}^{q}$
$\leq i^{\xi q}\sqrt{\mathbb{E}\left[\left(\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right)^{2q}\right]}\sqrt{\mathbb{E}\left[\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)\right\}^{2q}\right]}$
$=i^{(\xi-b-a/2)q}\sqrt{\mathbb{E}\left[i^{2qb}\left(\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right)^{2q}\right]}\sqrt{\mathbb{E}\left[i^{aq}\left\{\hat{f}_{i-1}\left(X_{i};Z_{i-1}\right)-f_{0}\left(X_{i}\right)\right\}^{2q}\right]}$
@@
$\mathbb{c}_{n}=\sqrt{\sum_{j=1}^{n-1}\left(\sum_{i=j+1}^{n}i^{\mathbb{c}}\right)^{2}}\leq c(n^{3/2+\mathbb{c}}\log n\vee\log n).$
So overall, we have
$\mathbb{P}\left(\left|\sum_{i=1}^{n}\mathbb{E}\left[u_{i}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}\right]\right|\geq c_{n}\right)$
$=$ $\mathbb{P}\left(\left|\sum_{i=1}^{n}\mathbb{E}\left[u_{i}\mid F^{i-1}\right]-\mathbb{E}\left[u_{i}\right]\right|/\mathbb{c}_{n}\geq c_{n}/\mathbb{c}_{n}\right)$